# Two transitions in complex eigenvalue statistics: Hermiticity and integrability breaking

### v1

@@ III Spectral statistics of complex eigenvalues
In this section, we introduce and compare two different random point processes of $N$ points $z_{1},\ldots,z_{N}$ in the complex plane (including the real line as a special case), namely the 2dCG and the $D$-dimensional Poisson point process, in addition to the class AI, which can be well approximated by the 2dCG with $\beta=1.4$. These processes are characterised by a joint distribution $\mathcal{P}_{N}(z_{1},\ldots,z_{N})$. We will recall exact and approximate analytical expressions, as well as numerically generated distributions for the local spacing between complex eigenvalues.
@@
The NN is given for each point $z_{i}$ by finding the NN in radial (Euclidean) distance as $s_{i}^{\textsc{nn}}=\min_{j\neq i}|z_{i}-z_{j}|$, being realised say by $j=i_{0}$  [^2] Such point may not be unique, but the event of finding two neighboring point at the very same distance will be of measure zero.. The NNN is then at distance $s_{i}^{\textsc{nnn}}=\min_{j\neq i,i_{0}}|z_{i}-z_{j}|$. When determining the respective distributions numerically, we first normalise the mean density by unfolding, as described in Sec. IV.1, and then determine the NN and NNN for all $i=1,\ldots,N$ to obtain a distribution. We will be careful to only take points inside the so-called bulk of the spectrum, that are sufficiently away from the edge of the spectrum—at order $O(1/\sqrt{N})$ for the random matrix ensembles we study. In all our examples, the spectral density has compact support  [^3] It is expected based on numerical simulations that the spacing distribution at the edge is different .
III.1 The 2-dimensional Coulomb gas and a surmise
@@
The first example of a point process we compare our data with is the 2dCG featuring a logarithmic interaction among the $N$ point charges, with a strength that can be interpreted as an inverse temperature $\beta$. Choosing a Gaussian confining potential, the joint distribution in the complex plane reads:
@@
The normalisation constant is given by the complex Selberg integral, which can be solved analytically only for $\beta=0$ or 2. Compared to the standard statistical mechanics notation, where the Gaussian term has a coefficient $\beta/2$, the positions of the point charges are rescaled as $z_{i}\to\sqrt{2/\beta}z_{i}$. This rescaling allows us to take the limit $\beta\to 0$ and thus recover the 2d Poisson case  [^4] The effect of this rescaling on the log term amounts to adding a constant $(\beta/2)\ln\sqrt{2/\beta}$, which in turn becomes a multiplicative factor, $\exp[\frac{\beta N(N-1)}{4}\ln\sqrt{\beta/2}]$, in the joint distribution. In the limit $\beta\to 0$, this factor goes to 1. In addition, the three different local symmetry classes of non-Hermitian random matrices can be obtained: namely, Eq. (6) with $\beta=2$ corresponds to the complex eigenvalue distribution of the complex Ginibre ensemble (class A), which can be solved analytically, cf. Appendix A; in turn, $\beta=1.4$ and $\beta=2.6$ give a very good approximation of the NN spacing distribution of class AI and class AII, respectively [akemann2022spacing]. Below, we will show that this approximation extends to NNN, too. To date, no analytic results for the local NN and NNN spacing are known apart for $\beta=0,2$  [^5] For $N=2$ the spacing distribution is known for AI and AII [hamazaki2020universality, jaiswal2019universality], which are however not a good approximation for the large-$N$ limit as was discussed in [akemann2022spacing]. The numerically determined spacings in the 2dCG are shown in Fig. 2.
Figure 2: The numerically generated NN (left) and NNN (right) spacing distributions of the 2dCG for $\beta=0-2$, increasing in steps $0.2$ here for illustration. This also includes analytic formulas for 2d Poisson with $\beta=0$, (9) and (10) with the lowest maxima, and random matrix class A with $\beta=2$, (20) and (LABEL:ANNN), with highest maxima.
@@
Figure 3: The NN (9) (left) and NNN (10) (right) Poisson spacing distribution, respectively, with dimension $1\leq D\leq 2$ varying in steps $0.1$. Notice that in both cases the maximum moves from right to left when decreasing dimension $D$.
III.2 The D-dimensional Poisson Point Process
@@
At $D=1$, it recovers the familiar exponential $P^{\textsc{nn}}_{\textrm{Pois},1}(s)=e^{-s}$, whereas at $D=2$, it is Gaussian with a linear prefactor stemming from the area measure, see also [sakhr2006wigner]. In turn, the NNN spacing distribution can be derived by integrating over the joint distribution of NN and NNN, provided in e.g. [sa2020complex], to find
@@
Since both distributions (9) and (10) are analytic in $D$, we can continue them to arbitrary $D>0$, as illustrated in Fig. 3. Such a Poisson distribution with an effective dimension $21<D<$ 2 was observed e.g. on forest patches, cf. [Akemann_2024].
III.3 Non-Hermitian random matrix class AI
@@
It was found in [akemann2022spacing] that the NN spacing of class AI distribution can be very well fitted by a 2dCG at $\beta=1.4$, which is also numerically generated, see the next subsection III.1. In this work, we also compare to the NNN spacing of AI, at the same value of $\beta=1.4$, and find an excellent agreement, see Fig. 4 right. The quality of the best fit for $\beta$ (marked in bold) is shown in standard deviations $\sigma$ (17) and Kolmogorov-Smirnov distances $d$ (18) in Table 1, to be introduced in Subsection IV.3. The results for class A are presented in Table 2.
@@
Figure 4: The NN (left) and NNN (right) AI spacing distribution generated numerically (histograms) for 500 matrices of size $N=5000$. Also shown are comparisons to the NN and NNN spacing distribution of the 2dCG, with fitted values $\beta=1.3$ (red dashed) and $\beta=1.4$ (blue full curve). The comparison with NNN is new here.
@@ IV Methods
IV.1 Unfolding complex spectra
@@
$s_{i}^{\prime\textsc{nn}}$ $=|z_{i}-z_{i}^{\textsc{nn}}|\sqrt{\bar{\rho}(x_{i},y_{i})}$ (15a)
@@
IV.2 Complex spacing ratios
@@
IV.3 Comparison of numerical data
@@
V Results for the spin chain data
We now characterize the spectral statistics of the spin chain with dissipative disorder described by the non-Hermitian Hamiltonian (1). We do so by comparing its distributions with that of the 2d spectral statistics introduced in Sec. III, using the methods presented in Sec. IV. We focus on the NN and NNN spacing distributions for values of the disorder strength from $\gamma=0-20$, fitting the numerical distribution to either the 2dCG distributions and finding the corresponding $\beta$, or to a $D$-dimensional Poisson point process with $1\leq D\leq 2$. For $L=16$, we present NN, NNN, and complex spacing ratio data for over 450 disorder realizations, while for $L=12,14$, we show this data for 1000 disorder realizations. This section finishes by a qualitative discussion of the complex spacing ratios [sa2020complex], to complete the picture.
We expect the spectral statistics of our model to behave similarly to those of a XXZ spin-chain Hamiltonian with real disorder, which exhibits a crossover of the spectral statistics from chaos to integrability (for non-zero disorder strength) as the disorder strength is increased. In such a case, the NN spectral statistics changes from random matrix statistics approximated by the Wigner surmise $\propto se^{-as^{2}}$ for the GOE in this model to 1d Poisson, $\propto e^{-s}$. In 1d, there is universal level repulsion between the eigenvalues only in the first, chaotic regime, while in the second, 1d Poisson displays uncorrelated eigenvalues—a universal signature of generic integrable (deterministic) systems.
In the non-Hermitian (2d) case with imaginary disorder, where the eigenvalues lie in the complex plane, we expect the spectral statistics to exhibit a crossover between random matrix statistics of class AI in our case to that of the 2d Poisson point process—which has a linear repulsion from the area measure. We follow the proposal [akemann2019universal] to fit this crossover using the spectral statistics of a 2d Coulomb gas with some value for $\beta\geq 0$, which includes 2d Poisson at $\beta=0$ as a special case and approximates the AI class for $\beta\approx 1.3-1.4$. However, the crossover has more structure than initially expected, showing two different crossovers, as we summarise below and detail in subsequent subsections.
$\gamma=0.0$              $\gamma=0.05$               $\gamma=0.1$              $\gamma=0.15$               $\gamma=0.2$              $\gamma=0.25$
NN Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
NNN
Eigenvalues
On a technical level, our spin chain data need to be drawn away from the edges of the eigenvalue distributions, where the statistics may be different or not even universal, as discussed at the beginning of Section III. To do so, we focus on the bulk given by the denser inner part of the spectrum. We systematically select only that part of the spectrum for which the eigenvalue density is not less than some fraction of the maximal density. For $L=16$ spins with $\gamma=0.75-20$, we use $1/2$ of the maximal density; for $L=14$ and $L=12$, we use a $1/3$ of the maximal density as a cutoff for the eigenvalues to be included in the analysis. For $L=16$ spins with $\gamma<0.75$ we choose a cutoff equal or smaller to $1/4$ the maximal density because, as we discuss below, the spectrum is very peaked around a line parallel to the real line.
The 2d spectral statistics described in Section III has the following features. The NN distribution of the 2dCG has its leftmost and lowest maximum at $\beta=0$, see Fig. 2 left (blue curve). Increasing $\beta$ leads to a monotonous increase of the maximum in height and in position to the right (blue to red). The random matrix class AI expected at maximal chaos is the one with the lowest maximum of the three distinct random matrix classes (see Fig. 14), reached at approximately $\beta\approx 1.3-1.4$. As such, it is approached by the 2dCG monotonously from below, starting at $\beta=0$. What happens if the maximum of the NN spacing distribution of our spin chain data—which is normalised and has first moment of unity—is to the left of the 2d Poisson distribution (at $\beta=0$)? Obviously the 2dCG will no longer provide a good fit then. Only the $D$-dimensional Poisson distribution might provide a good fit then for $D<2$, see Fig. 3 left.
Our findings can be summarised by Fig. 6 and described as follows. In the small disorder regime, there is a crossover between 1d Poisson at $\gamma=0$—where our XXZ spin chain is integrable— towards 2d Poisson, that is well described by an increase in effective dimension $21\leq D<$ 2. We reach up to $D\approx 1.7$ at $\gamma=0.25$ for both NN and NNN spacing distributions. Therefore, Hermiticity is broken first on a smaller scale. This is summarised in Fig. 6 left and described in detail in Subsection V.1.
For intermediate to large values of $\gamma=0.5-20$, we see a rise from $\beta\approx 0.6$ at $\gamma=0.5$ to $\beta\approx 1.3-1.4$ at $\gamma=2$ where class AI is located, to which we also compare. We see that integrability is broken as we approach the fully chaotic behaviour characterised by random matrix class AI. For larger $\gamma$, the disorder term (3) leads to an alignment of spins and thus again to a deterministic behaviour, approaching 2d Poisson with $\beta=0$ around $\gamma\approx 8$, see Fig. 6 right. This is discussed in Subsection V.2, including the approach to class AI with system size $L$ in Subsection V.2.1. The asymptotic approach to $\beta=0$ between $\gamma=8-20$ is discussed in Subsection V.3, where we also use the surmise (7) for the finer scaling in the NN distribution. To complete the picture, we show and discuss results for the complex spacing ratios at various values of $\gamma$ in Subsection V.4.
V.1 Hermiticity breaking at small disorder $\gamma$:
from 1d to 2d Poisson
$\gamma=0.5$               $\gamma=1$               $\gamma=2$                $\gamma=3$               $\gamma=4$               $\gamma=5$                $\gamma=10$
Eigenvalues
Starting with $\gamma=0$ (left plots), the spectrum is real, the XXZ model integrable, and we obtain a very good agreement with 1d Poisson statistics for both NN and NNN as expected. Notice that at $\gamma=0$ the energies are doubly degenerate, which is why we only show spacings with $\mathcal{S}_{z}=0$ and parity $P=+1$ (for $L=16$, this sector is of dimension 6470). The plots for $\mathcal{S}_{z}=0$ with $P=-1$ (for $L=16$, this sector is of dimension 6400) look similar. For our smallest value of $\gamma$, we obtain a somewhat reasonable fit for the NN, with $D\approx 0.88$ consistent with $D=1$. In contrast, the distribution for the NNN does not yield a good fit at all, a feature that persists for $\gamma=0.1$. A possible explanation is that for these small values of $\gamma$ the spectrum is almost singular, being concentrated on a line parallel to the real line, with very few eigenvalues in the complex plane, see Fig. 8 below, and the unfolding procedure described in Section IV.1 for a 2d complex spectrum may then not be completely suitable. For $\gamma=0$, we used the 1d version of the unfolding method described in IV.1  [^9] We found that for 1d, $\Sigma=1.8\bar{s}$ gives good results..
From $\gamma=0.1$ both NN and NNN start to deviate from $D=1$, and the best fits in $D$ reach values up to $D\approx 1.7$ at $\gamma=0.25$. Comparing NN and NNN leads to consistent values within a small margin. We thus believe the fits to a one-parameter family of distributions is not a coincidence, but rather capture properties of an underlying effective Poisson interaction.
This leads to the interpretation that Hermiticity is broken first, while maintaining an (almost) integrable system, the characteristic of which is given by a Poisson distribution at intermediate dimension $21<D<$ 2. Increasing to the next higher value of $\gamma=0.5$, detailed in the next subsection V.2, the maximum of the NN crosses to the right of the 2d Poisson distribution (at $\beta=0$), which then fits the 2dCG. The fact that $D=2$ is not fully reached by increasing $\gamma$ here can be interpreted as a competition between Hermiticity and integrability breakings around the scale $\gamma=0.25-0.5$.
As an example for the crossover between one- and two-dimensional spectra, we show in Fig. 8 cuts through the spectral density perpendicular to the real axis at $x\approx 0$ as a function of imaginary part $y$. Noting the difference in scale, we see that between $\gamma=0.1$ and $0.2$ shown here, the spectrum slowly softens and starts spreading into the complex plane. The difference in abundance of eigenvalues with non-vanishing imaginary part may explain why the NNN distribution in Fig. 7 fits well only from a certain value of $\gamma=0.15$ onwards.
Table 3: List of best fit $\beta$ for the comparison of the NN spacing and NNN spacing of the spin chain with $L=16$ and $\gamma=0.5,0.75,1,2,\ldots 20$ to the 2dCG and the corresponding standard deviation in units of $10^{-2}$. Notice that apart from $\gamma=0.5$ the values obtained for $\beta$ from NN and NNN agree or differ by $0.1$ only.
V.2 Integrability breaking at intermediate to large disorder $\gamma$: from 2d Poisson to class AI and back
In this section, we discuss the transition from random matrix statistics class AI to 2d Poisson statistics, witnessed for intermediate ($\gamma=0.5$) to large ($\gamma=20$) disorder dissipation strength. The approach towards fully chaotic statistics represented by random matrix class AI when increasing system size $L=12,14$ and 16 is included in subsection V.2.1 below.
The smallest value of disorder for which a meaningful fit with a 2dCG at $\beta\geq 0$ can be made is $\gamma=0.5$, when, $\beta=0.6\,(0.4)$ for NN (NNN), see Fig. 9 left. At this disorder strength, both Hermiticity and integrability appear to be broken. The system is in between the integrable case ($\beta=0$) and the fully chaotic case ($\beta\approx 1.3-1.4$, equivalent to random matrix symmetry class AI). The latter is then attained at $\gamma=2$, where the maximal value for $\beta$ (and height of the maximum) is reached, see Figure 6 right. We compare here the NN and NNN spacings with the 2dCG rather than AI. This is because the AI class is very well approximated by the 2dCG at these values of $\beta$, as shown in Fig. 4 and Table 1. Furthermore, random matrix statistics is approached in $\beta$ from below, see also the next subsection V.2.1, and the $\beta$-value gives a measure of how close it is.
Increasing $\gamma$ further leads to a rapid decrease in the best fitted value for $\beta$, see also Table 3 below. From $\gamma\simeq 8-10$ onwards we consistently obtain $\beta=0$. This is because the disorder is so strong that the spins align with the disorder term and the statistics becomes again deterministic. The best fits for $\beta$ for NN and NNN and corresponding values, including standard deviations, are collected in Table 3. In this section we only fit to $\beta$ in the 2dCG where the step width is 0.1, but both NN and NNN distribution are available. The next section V.3 presents a much finer fit in $\beta$ for $\gamma\geq 8$, possible for the NN only thanks to the surmise (7), which is a very good approximation so close to $\beta=0$. The approach to zero can also be seen well in the summary Fig. 6 right. Similar findings for Hermitian Hamiltonians comparing to 1d Poisson and corresponding classical random matrices, using both NN and NNN spacings, were made very recently [kundu2024beyond].
$L=12$                $L=14$               $L=16$
NN
NNN
Figure 10: Comparison of the XXZ with dissipative order $\gamma=2$ in the $\mathcal{S}_{z}=0$ magnetization sector with increasing number of spins $L=12,14,16$ (from left to right, histograms) and the symmetry class AI of complex symmetric matrices (blue curve) of size $N=5000$. The NN distribution is given in the first row, whereas the NNN spacing is shown in the second row. Additionally, we give the standard deviation between the two plotted distributions below each plot. We observe a decrease of the standard deviation for growing $L$, as expected.
$L=12$                $L=14$               $L=16$
NN
$\sigma=1.8\cdot 10^{-2}$          $\sigma=0.6\cdot 10^{-2}$         $\sigma=0.8\cdot 10^{-2}$
NNN
$\sigma=1.7\cdot 10^{-2}$          $\sigma=0.5\cdot 10^{-2}$          $\sigma=1.3\cdot 10^{-2}$
Table 4: List of best fit $\beta$ for the comparison of the nearest-neighbour (NN) spacing of the spin chain with $L=16$ and $\gamma=8,\ldots 20$ to the NN 2dCG surmise (7) and the corresponding standard deviation in units of $10^{-2}$.
$\boldsymbol{\gamma}$ 8 9 10 11 12 13 14 15 16 17 18 19 20
$\sigma$ 0.36 0.35 0.33 0.35 0.33 0.33 0.31 0.3 0.27 0.24 0.24 0.21 0.2
V.2.1 Dependence on system size $L$
The main data presented here are always for the longest spin chain, with $L=16$. This corresponds to a Hilbert space dimension of $12870$, that allows reaching a fully chaotic behaviour. Here, we also present data for system sizes $L=14$ and $L=12$, with respective dimensions of Hilbert space $3432$ and $924$, at the values $\gamma=2$ and $\gamma=20$. While for all three system sizes the 2d Poisson limit is reached fully at large $\gamma=20$, see Fig. 11, this is not the case for the limit towards random matrix statistics of class AI, see Fig. 10.
size $L$ 12 14 16
NN $\beta$ 0.9 1.2 1.3
$\sigma$ 1.3 0.6 0.7
NNN $\beta$ 0.8 1.1 1.3
$\sigma$ 3.5 1.3 1.0
V.3 Asymptotic approach to 2d Poisson for large disorder $\gamma$
V.4 Comparison to complex spacing ratio discussion
The probability density of the complex spacing ratios, as seen in Figure 13, contains interesting patterns especially in the chaotic regime. The lower density around the origin and around $r=1$ are due to level repulsion. It is not likely for an eigenvalue and its NN to be close together, thus the density is lower at $r=0$; and an eigenvalue’s NN and NNN would most likely lie on either side of it, making the density lower at $r=1$. The ‘bow and arrow’ feature seen for small $\gamma$ is a manifestation of an underlying structure of the complex eigenvalues. As observed and explained in Appendix D of [Li:2024uzg], this feature is a result of an eigenvalue’s NN and NNN being on the real line and as complex conjugate to it, or vice versa. Section II.2 showed that the Hamiltonian is pseudo-Hermitian and thus its spectrum is symmetric to reflection across a line parallel to the real axis. In the complex spacing ratios, this symmetry manifests itself as reflection symmetry across the real axis. Additionally, the excess of ratios along the real line is a result of real eigenvalues having real NN and NNN.
$\gamma=0.05$            $\gamma=0.1$           $\gamma=0.15$              $\gamma=0.5$             $\gamma=2$             $\gamma=5$              $\gamma=10$
CSR Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption

### v2

@@ IV Spectral statistics of complex eigenvalues
In this section and in the next we describe the tools we have been using to analyze our results in the previous section. In particular, we introduce and compare two different random point processes of $N$ points $z_{1},\ldots,z_{N}$ in the complex plane (including the real line as a special case), namely the 2dCG and the $D$-dimensional Poisson point process, in addition to the class AI, which can be well approximated by the 2dCG with $\beta=1.4$ [akemann2022spacing]. These processes are characterised by a joint distribution $\mathcal{P}_{N}(z_{1},\ldots,z_{N})$. We will recall exact and approximate analytical expressions, as well as numerically generated distributions for the local spacing between complex eigenvalues.
@@
The NN is given for each point $z_{i}$ by finding the NN in radial (Euclidean) distance as $s_{i}^{\textsc{nn}}=\min_{j\neq i}|z_{i}-z_{j}|$, being realised say by $j=i_{0}$  [^3] Such point may not be unique, but the event of finding two neighboring point at the very same distance will be of measure zero.. The NNN is then at distance $s_{i}^{\textsc{nnn}}=\min_{j\neq i,i_{0}}|z_{i}-z_{j}|$. When determining the respective distributions numerically, we first normalise the mean density by unfolding, as described in Sec. V.1, and then determine the NN and NNN for all $i=1,\ldots,N$ to obtain a distribution. We will be careful to only take points inside the so-called bulk of the spectrum, that are sufficiently away from the edge of the spectrum—at order $O(1/\sqrt{N})$ for the random matrix ensembles we study. In all our examples, the spectral density has compact support  [^4] It is expected based on numerical simulations that the spacing distribution at the edge is different .
IV.1 The 2-dimensional Coulomb gas and a surmise
@@
The first example of a point process we compare our data with is the 2dCG, featuring a logarithmic interaction among the $N$ point charges, with a strength that can be interpreted as an inverse temperature $\beta$. Choosing a Gaussian confining potential, the joint distribution in the complex plane reads:
@@
The normalisation constant is given by the complex Selberg integral, which can be solved analytically only for $\beta=0$ or 2. Compared to the standard statistical mechanics notation, where the Gaussian term has a coefficient $\beta/2$, the positions of the point charges are rescaled as $z_{i}\to\sqrt{2/\beta}z_{i}$. This rescaling allows us to take the limit $\beta\to 0$ and thus to recover the 2d Poisson case  [^5] The effect of this rescaling on the log term amounts to adding a constant $(\beta/2)\ln\sqrt{2/\beta}$, which in turn becomes a multiplicative factor, $\exp[\frac{\beta N(N-1)}{4}\ln\sqrt{\beta/2}]$, in the joint distribution. In the limit $\beta\to 0$, this factor goes to 1. In addition, the three different local symmetry classes of non-Hermitian random matrices can be obtained: namely, Eq. (6) with $\beta=2$ corresponds to the complex eigenvalue distribution of the complex Ginibre ensemble (class A), which can be solved analytically, cf. Appendix A; in turn, $\beta=1.4$ and $\beta=2.6$ give a very good approximation of the NN spacing distribution of class AI and class AII, respectively [akemann2022spacing]. Below, we will show that this approximation extends to NNN, too. To date, no analytic results for the local NN and NNN spacing are known apart for $\beta=0,2$  [^6] For $N=2$ the spacing distribution is known for AI and AII [hamazaki2020universality, jaiswal2019universality], which are however not a good approximation for the large-$N$ limit as was discussed in [akemann2022spacing]. The numerically determined spacings in the 2dCG are shown in Fig. 2(b).
@@
IV.2 The D-dimensional Poisson Point Process
@@
At $D=1$, it recovers the familiar exponential distribution $P^{\textsc{nn}}_{\textrm{Pois},1}(s)=e^{-s}$, whereas at $D=2$, it is Gaussian with a linear prefactor stemming from the area measure, see also [sakhr2006wigner]. In turn, the NNN spacing distribution can be derived by integrating over the joint distribution of NN and NNN, provided in e.g. [sa2020complex], to find
@@
Since both distributions (9) and (10) are analytic in $D$, we can continue them to arbitrary $D>0$, as illustrated in Fig. 2(a). Such a Poisson distribution with an effective dimension $21<D<$ 2 was observed e.g. on forest patches, cf. [Akemann_2024].
IV.3 Non-Hermitian random matrix class AI
@@
It was found in [akemann2022spacing] that the NN spacing of class AI distribution can be very well fitted by a 2dCG at $\beta=1.4$, which is also numerically generated, see the next subsection IV.1. In this work, we also compare to the NNN spacing of AI, at the same value of $\beta=1.4$, and find an excellent agreement, see Fig. 11 right. The quality of the best fit for $\beta$ (marked in bold) is shown in standard deviations $\sigma$ (17) and Kolmogorov-Smirnov distances $d$ (18) in Table 4, to be introduced in Subsection V.3. The results for class A are presented in Table 5.
@@
Figure 11: The NN (left) and NNN (right) AI spacing distribution generated numerically (histograms) for 500 random matrices of size $N=5000$. Also shown are comparisons to the NN and NNN spacing distribution of the 2dCG, with fitted values $\beta=1.3$ (red dashed) and $\beta=1.4$ (blue full curve). The comparison with NNN is new here.
@@ V Methods
V.1 Unfolding complex spectra
@@
$s_{i}^{\prime\textsc{nn}}$ $=|z_{i}-z_{i}^{\textsc{nn}}|\sqrt{\bar{\rho}(x_{i},y_{i})},$ (15a)
@@
V.2 Complex spacing ratios
@@
V.3 Comparison of numerical data